OpenStudy (anonymous):

solve x - 8 greater than or equal to -5(x + 3) write a solution using builder notation; show work

OpenStudy (anonymous):

x - 8 >= -5(x+3) x - 8 >= -5x - 15 x + 5x >= -15 + 8 6x >= -7 x >= -7/6

OpenStudy (anonymous):

ok thanks